4. A) What is the rate law that corresponds to the data shown for the reaction 2A B-C? Exp Initial [A] IniiB] Inii rate 0.015 0.030 0.060 0.060 0.022 0.044 0.044 0.066 0.125 0.500 0.500 1.125 4 b) Determine the half-life of the reaction.

c) What is the order in the reaction; zero order, 1st order, or 2nd order.

Please show all work, please don't just say that

Step1 From 2nd and 3rd observations [A] has doubled; [B] has remained the same and rate hs also remained the same. Thus order with respect to A is 0.

Step2 From 3rd and 4th observations [A] has remained same ;[B] has become 1.5 times;
the rate has become 2.25 times. Thus order with respect to B is 2.
Answer #1

when concentration of A changed alone rate does not change so with respect to A reaction is zero order

when concentration of B doubled rate increased 4 times so with respect to B reaction is second order

so rate law as follows

rate = K[B]2

so overall order = 2

for second order

t1/2 = 1 / K[B0]

so first we must find K for the reaction

K = rate / [B]2

K = 0.125 / [0.022]2 = 0.125 / 0.000484

K = 2.58 x 102 M-1s-1

t1/2 = 1 / 2.58 x 102 x 0.022

t1/2 = 1 / 5.676

t1/2 = 0.176 s

what you have shown in step 1 and step 2 is correct only